What is the mass of water which contains 50% heavy water (D2O) ?

Grade:10

2 Answers

Arun
6 years ago
Dear simran
 
For 1 mole water the mass is = 18 g
 
For 1 mole D2O the mass is = 20 g
 
If both are present in 50/50 amount
 
mass = (18 + 20)/2 = 38/2 = 19 g
 
Regards
Arun (askIITiams forum expert)
Subhash Kumar Mehta
6 years ago
50 percent of heavy water and normal water means average mass of normal water and heavy water =18+20/2»=19 gram »» i think you would understood it well dear #Simran Regards»#Subhash_Kumar_Mehta